# Learning Autonomous Surgical Irrigation and Suction with the da Vinci Research Kit Using Reinforcement Learning

### v1

@@ Abstract
The irrigation-suction process is a common procedure to rinse and clean up the surgical field in minimally invasive surgery (MIS). In this process, surgeons first irrigate liquid, typically saline, into the surgical scene for rinsing and diluting the contaminant, and then suction the liquid out of the surgical field. While recent advances have shown promising results in the application of reinforcement learning (RL) for automating surgical subtasks, fewer studies have explored the automation of fluid-related tasks. In this work, we explore the automation of both steps in the irrigation-suction procedure and train two vision-based RL agents to complete irrigation and suction autonomously. To achieve this, a platform is developed for creating simulated surgical robot learning environments and for training agents, and two simulated learning environments are built for irrigation and suction with visually plausible fluid rendering capabilities. With techniques such as domain randomization (DR) and carefully designed reward functions, two agents are trained in the simulator and transferred to the real world. Individual evaluations of both agents show satisfactory real-world results. With an initial amount of around 5 grams of contaminants, the irrigation agent ultimately achieved an average of 2.21 grams remaining after a manual suction. As a comparison, fully manual operation by a human results in 1.90 grams remaining. The suction agent achieved 2.64 and 2.24 grams of liquid remaining across two trial groups with more than 20 and 30 grams of initial liquid in the container. Fully autonomous irrigation-suction trials reduce the contaminant in the container from around 5 grams to an average of 2.42 grams, although yielding a higher total weight remaining (4.40) due to residual liquid not suctioned. Further information about the project is available at https://tbs-ualberta.github.io/CRESSim/.
@@ I Introduction
Agrowing interest has been shown in recent years to achieve autonomous or semi-autonomous subtask execution in surgeries using surgical robotic systems. These include tissue cutting [1, 2, 3], tissue manipulation [4, 5, 6, 7], suturing [8, 9, 10], and others, with the overall goal of reducing the workload and enhancing surgeons’ skills when performing specific subtasks, achieving what is termed augmented dexterity—the enhancement of a surgeon’s capabilities through robotic assistance under supervision [11]. Increasing numbers of these studies are utilizing machine learning approaches, particularly robot learning techniques such as reinforcement learning (RL) and imitation learning (IL), due to their generalizability and reduced manual effort when compared with traditional motion planning algorithms, thanks to their data-driven nature.
While these studies show promising results, many rely on manually extracted feature vectors from images as input to the policy, such as tooltip positions and key point locations extracted from markers placed on the surgical instrument or tissue. Although technically feasible, this can be less robust in realistic surgical scenarios, such as when the lighting condition changes and when smoke is present due to energy-based operations. Therefore, relying directly on the raw visual observation captured from the endoscopic camera as input to the policy is usually more desirable, due to the elimination of manual feature extraction and the potential of obtaining an end-to-end policy that takes the sensory input (e.g., image and sensor reading) and directly produces the desired actions.
@@
However, one major challenge when applying sim-to-real approaches to surgical task automation is the need to simulate realistic surgical scenes and operations. Unlike simulators for everyday and industrial robotic tasks, which typically involve mostly rigid objects governed by basic principles of mechanics, surgical simulations require a higher level of complexity to accurately model soft tissues, fluids, and dynamic interactions such as cutting and cauterization. The simulation of such physical models and behaviors is already a considerable challenge, but rendering visually realistic images in the simulator to support the training of vision-based agents adds an additional layer of complexity.
@@
These complexities are particularly evident in the underexplored area of surgical irrigation and suction. During surgery, blood, debris, and other contaminants are frequently present as a result of procedures like cutting, cauterization, and tissue manipulation. To maintain a clear surgical field, surgeons may follow a two-step process: first, they irrigate the area with a sterile solution, typically saline, to rinse or dilute the contaminants; then, suction is used to remove the fluid along with any remaining debris. This process is usually carried out using an irrigation-suction instrument, which allows the surgeon to both introduce and remove fluids. However, this involves a considerable amount of time and effort and leads to intermittent interruptions of the primary surgical procedure. Therefore, automating this process can help reduce the workload of surgeons.
@@
Developing simulated learning environments for the task is challenging, as it involves fluid dynamics and the corresponding interactions for irrigating and suctioning fluids. Fluid mixing should also be considered, as the irrigated solution will be mixed with the existing contaminants. Additionally, developing a vision-based policy involves a greater degree of complexity since the rendering of fluids while mixing them is also required to synthesize visually plausible images.
@@
Existing studies that attempt to automate this task primarily focus on the suction phase, particularly blood suction [13, 14, 15, 16, 17]. In [14], the authors employ optical flow to track blood flow in real-time and generate a suction trajectory using a computational motion planner to efficiently remove blood while moving upstream toward the bleeding point. Huang et al. implemented a differentiable position-based fluids (PBF) model of the blood and used model predictive control (MPC) to generate a suction trajectory. In our previous work, a simulated blood suction environment is built based on PBF and is used to train an RL agent that takes a binary image mask of the blood region as input, and outputs the 2D motion of the suction tool [16]. However, manual feature extraction is still needed to obtain a binary mask representing the area of the blood. To the best of the authors’ knowledge, none of the existing studies have developed an agent relying on raw RGB images without manual feature extraction, and the automation of irrigation has not been explored in the literature.
@@
In this work, we develop vision-based agents that autonomously complete both the two steps during surgical irrigation and suction. To achieve this, we build a novel surgical robot learning platform for the da Vinci Research Kit (dVRK) that integrates robot learning capabilities and visually realistic fluid simulations and rendering. We design irrigation and suction learning environments in the simulator that resemble the real-world setup. Vision-based agents are trained in the simulator with domain-randomized parameters and carefully designed rewards, learning curricula, and human demonstrations, and are then transferred to the real world. The main contributions of this work are as follows:
  • We propose CRESSim-ML, a surgical robot learning platform for the dVRK, built on our previous work (CRESSim) [18]. The platform is developed with Unity, ML-Agents, and PhysX 5 SDK, and incorporates the actual dVRK Master Tool Manipulator (MTM) for collecting human demonstrations through teleoperation.
  • Based on CRESSim-ML, simulated learning environments are created for irrigation and suction using the Patient Side Manipulator (PSM) from the dVRK and the EndoWrist Suction/Irrigator. Screen-space fluid rendering is implemented and adapted to simulate realistic visual effects including fluid mixing.
  • With carefully designed reward functions, DR, curriculum learning (CL), and IL, two vision-based agents for irrigation and suction are trained in the simulator for controlling the robot in joint space to complete the tasks.
@@ II-A Surgical Subtask Automation and Augmented Dexterity
A large number of recent studies aim to achieve subtask automation in surgeries using surgical robotic systems, with a long-term goal of increasing the level of autonomy in surgeries, similar to autonomous vehicles. The concept where surgical subtasks are controlled by a robot under surgeons’ supervision is recently referred to as augmented dexterity [11]. Examples of these studies include tissue cutting [1, 2, 3], tissue manipulation [5, 6, 7], suturing [8, 9, 10], blood suction [13, 14, 15, 16, 17], and vessel manipulation [19]. Numerous studies have also focused on the navigation and control of endoscopes [20, 21, 22]. Many of these studies leverage RL approaches [1, 2, 3, 5, 6, 7, 9, 10, 8, 16]. However, applying RL to irrigation and suction in surgery has been less investigated, largely due to the complexity of simulating a training environment that involves fluid manipulation.
@@ II-B Surgical Simulation and Sim-to-Real Transfer
Simulating surgical scenes is challenging compared with daily tasks, due to the multiple types of objects and manipulations involved, such as soft bodies (e.g., soft tissue), fluids (e.g., blood and other body fluids), burning, and cutting. Nevertheless, its applications have been important in two major domains. The traditional usage is for training surgeons, where trainees operate in a simulated surgical environment to improve their skills. Examples of such systems include da Vinci SimNow111www.intuitive.com/en-us/products-and-services/da-vinci/learning/simnow, VirtaMed LaproS222www.virtamed.com/en/products-and-solutions/simulators/laparos, and iMSTK333www.imstk.org.
A more emerging application of surgical simulation is for training autonomous agents using robot learning methods, where agents are trained in the simulator and transferred to the real world. Studies of this type include [4, 7, 9, 10, 12, 16]. While it is ideal to use commercial simulators such as the da Vinci SimNow for building training environments for robot learning, it is challenging in practice due to the proprietary nature of these software systems. Many recent studies have proposed open-source simulation environments for surgical robot learning, including dVRL [23], AMBF-RL [24], UnityFlexML [4], SurRoL [25, 26], LapGym [27], and Surgical Gym [28]. However, they are generally inferior to the commercial ones in terms of the types of objects that they can simulate. For instance, AMBF-RL is mainly used for simulating rigid bodies and some soft bodies, without the capability for simulating fluids. Nevertheless, achieving sim-to-real transfer for irrigation-suction requires the simulation of visually plausible fluids.
@@ II-C Fluid Simulation and Manipulation
This work is also related to recent advancements in robotic manipulation involving fluids, as well as the development of fluid simulation environments for robot learning [29, 30, 31]. In [29], the Navier–Stokes equation is used for simulating 2D fluid flow, and an RL model is trained to learn to manipulate a rigid object using a water jet. Babaians et al. used PBF for simulating fluids [30], which is used to train a robot to pour a glass of liquid. Authors of [31] considered a number of fluid manipulation tasks, such as fluid mixing and drawing latte art. Most of these studies focus on daily tasks such as pouring and beverage mixing, which is substantially different from surgical irrigation or suction in terms of task dynamics.
@@ III-A1 Fluid Simulation With Color Diffusion
To simulate irrigation and suction, fluid behavior must be simulated and rendered with visual plausibility. PhysX 5 SDK provides the feature of simulating fluid particles on GPU using position-based dynamics (PBD), and more specifically, position-based fluids (PBF) [32].
@@
As we would like to mix two types of fluids with different colors when simulating irrigation, the simulation should include color mixing and changes. It is therefore straightforward to label each particle with a color and diffuse the colors between particles based on their spatial proximity. To achieve this, each particle is assigned to a cell in a 3D grid with uniform discrete cells based on its position, also known as spatial hashing. For each particle, all particles in a neighboring region are evaluated and the weighted average of the particle colors is computed to update the particle’s color. Therefore, for each particle $i$ with color $\mathbf{c}_{i}$,
@@ III-A2 Fluid Rendering
Figure 3: Example fluid simulation and rendering scene. Fluids generated from two fluid sources of different colors are mixed together.
@@
However, rendering should account for the different colors of each particle in our case. To achieve this, an additional color extraction pass is added to the regular screen-space rendering scheme after the surface depth texture is extracted, which simply extracts the colors of particles that are closest to the fluid surface. This process is similar to [35], where the authors propose using a separate pass to extract the colors based on the thickness of the fluid. The difference is that we do not consider any particle colors apart from the surface ones. This eliminates the need for a thickness pass to extract the fluid thickness and is a valid simplification without losing much visual plausibility thanks to the previously discussed color diffusion scheme.
We use anisotropy ellipsoids discussed in [36] when reconstructing the fluid surface and a narrow range filter [37] to smooth it. An example fluid simulation and rendering scene is shown in Fig. 3. The overall fluid rendering steps are shown in Fig. 4.
@@ III-B A Surgical Robot Learning Framework for the dVRK
In our previous work [18], we have built a general surgical simulation platform (CRESSim) by leveraging PhysX 5 and Unity, where the dVRK PSM robot is simulated and the real-world dVRK MTM robot is used for teleoperating the simulated PSM. The PSM robot together with the EndoWrist One Suction/Irrigator is simulated as a kinematic tree with multiple articulation joints, as shown in Fig. 5. It is worth noting that the Suction/Irrigator has only two degrees of freedom (DoF), making the entire PSM a 5-DoF robot.
@@
We are able to further implement a robot learning framework by incorporating the Unity Machine Learning Agents Toolkit (ML-Agents) [38] into the existing software. Human demonstrations for completing simulated tasks can then be collected while using the real-world MTM to teleoperate the PSM in the simulated learning scene, as shown in Fig 6. The procedures for collecting demonstrations will be further discussed in Section IV-B.
Figure 6: A surgical robot learning platform for the dVRK with Unity, ML-Agents, and PhysX 5.
Based on this platform, two simulated learning scenes for irrigation and suction are developed, which will be discussed in the following section. While we only consider irrigation and suction in this work, the platform can be generally used for building various surgical robot learning tasks for the dVRK, including those that use different instruments such as the Large Needle Driver. Our previous work has shown its capability to simulate a number of surgical tasks, such as tissue cutting and manipulation [18]. As the proposed platform combines CRESSim with ML-Agents, we refer to it as CRESSim-ML.

### v2

@@ Abstract
The irrigation-suction process is a common procedure to rinse and clean up the surgical field in minimally invasive surgery (MIS). In this process, surgeons first irrigate liquid, typically saline, into the surgical scene for rinsing and diluting the contaminant, and then suction the liquid out of the surgical field. While recent advances have shown promising results in the application of reinforcement learning (RL) for automating surgical subtasks, fewer studies have explored the automation of fluid-related tasks. In this work, we explore the automation of both steps in the irrigation-suction procedure and train two vision-based RL agents to complete irrigation and suction autonomously. To achieve this, a platform is developed for creating simulated surgical robot learning environments and for training agents, and two simulated learning environments are built for irrigation and suction with visually plausible fluid rendering capabilities. With techniques such as domain randomization (DR) and imitation learning, two agents are trained in the simulator and transferred to the real world. Individual evaluations of both agents show satisfactory real-world results. With an initial amount of around 5 grams of contaminants, the irrigation agent ultimately achieved an average of 2.21 grams remaining after a manual suction. As a comparison, fully manual operation by a human results in 1.90 grams remaining. The suction agent achieved 2.64 and 2.24 grams of liquid remaining across two trial groups with more than 20 and 30 grams of initial liquid in the container. Fully autonomous irrigation-suction trials reduce the contaminant in the container from around 5 grams to an average of 2.42 grams, although yielding a higher total weight remaining (4.40) due to residual liquid not suctioned. Further information about the project is available at https://tbs-ualberta.github.io/CRESSim/.
@@ I Introduction
Agrowing interest has been shown in recent years in achieving autonomous or semi-autonomous subtask execution in surgeries using surgical robotic systems. These include tissue cutting [1, 2, 3], tissue manipulation [4, 5, 6, 7], suturing [8, 9, 10], and others, with the overall goal of reducing the workload and enhancing surgeons’ skills when performing specific subtasks, achieving what is termed augmented dexterity—the enhancement of a surgeon’s capabilities through robotic assistance under supervision [11]. Increasing numbers of these studies are utilizing machine learning approaches, particularly robot learning techniques such as reinforcement learning (RL) and imitation learning (IL), due to their generalizability and reduced manual effort when compared with traditional motion planning algorithms, thanks to their data-driven nature.
While these studies show promising results, many rely on manually extracted feature vectors from images as input to the policy, such as tooltip positions and key point locations extracted from markers placed on the surgical instrument or tissue. Although technically feasible, this can be less robust in realistic surgical scenarios, such as when the lighting conditions change and when smoke is present due to energy-based operations. Therefore, relying directly on the raw visual observation captured from the endoscopic camera as input to the policy is usually more desirable, due to the elimination of manual feature extraction and the potential of obtaining an end-to-end policy that takes the sensory input (e.g., image and sensor reading) and directly produces the desired actions.
@@
However, one major challenge when applying sim-to-real approaches to surgical task automation is the need to simulate realistic surgical scenes and operations. Unlike simulators for everyday and industrial robotic tasks, which typically involve mostly rigid objects governed by basic principles of mechanics, surgical simulations require a higher level of complexity to accurately model soft tissues, fluids, and dynamic interactions such as cutting and cauterization. The simulation of such physical models and behaviors is already a considerable challenge involving computational solid and fluid mechanics, but rendering visually realistic images for fluid in the simulator to support the training of vision-based agents adds an additional layer of complexity.
@@
These complexities are particularly evident in the underexplored area of surgical irrigation and suction. During surgery, blood, debris, and other contaminants are frequently present as a result of procedures like cutting, cauterization, and tissue manipulation [14]. To maintain a clear surgical field, surgeons may follow a two-step process: first, they irrigate the area with a sterile solution, typically saline, to rinse or dilute the contaminants; then, suction is used to remove the fluid along with any remaining debris. This process is usually carried out using an irrigation-suction instrument in robotic surgery, which allows the surgeon to both introduce and remove fluids. However, this usually involves a considerable amount of time and effort [15], and automating this process may help reduce the workload of surgeons.
@@
Developing simulated learning environments for the task is challenging, as it involves fluid dynamics, which is more computationally complex than rigid body dynamics, and the corresponding interactions for irrigating and suctioning fluids. Fluid mixing should also be considered, as the irrigated solution will be mixed with the existing contaminants. Additionally, developing a vision-based policy involves a greater degree of complexity since the rendering of fluids while mixing them is also required to synthesize visually plausible images.
@@
Existing studies that attempt to automate this task primarily focus on the suction phase, particularly blood suction [16, 17, 18, 19, 20]. In [17], the authors employ optical flow to track blood flow in real-time and generate a suction trajectory using a computational motion planner to efficiently remove blood while moving upstream toward the bleeding point. Huang et al. implemented a differentiable position-based fluids (PBF) model of the blood and used model predictive control (MPC) to generate a suction trajectory. In [19], a simulated blood suction environment is built based on PBF and is used to train an RL agent that takes a binary image mask of the blood region as input, and outputs the 2D motion of the suction tool. However, manual feature extraction is still needed to obtain a binary mask representing the area of the blood. To the best of the authors’ knowledge, none of the existing studies have developed an agent relying on raw RGB images without manual feature extraction, and the automation of irrigation has not been explored in the literature.
@@
In this work, we develop vision-based agents that autonomously complete both steps during surgical irrigation and suction. To achieve this, we build a novel surgical robot learning platform for the da Vinci Research Kit (dVRK) [21] that integrates robot learning capabilities and visually realistic fluid simulations and rendering. We design irrigation and suction learning environments in the simulator that resemble the real-world setup. Vision-based agents are trained in the simulator with domain-randomized parameters and carefully designed rewards, learning curricula, and human demonstrations, and are then transferred to the real world. This approach is similar to the one used in [13], where a vision-based rope-cutting model is obtained in a sim-to-real manner. The main contributions of this work are as follows:
  • We propose CRESSim-ML, a surgical robot learning platform for the dVRK. Built on our previous work (CRESSim) [22], the platform additionally incorporates parallel training environment capabilities, and allows collecting human demonstrations through teleoperation using the actual dVRK Master Tool Manipulator (MTM).
  • Based on CRESSim-ML and Unity’s ML-Agents, simulated learning environments are created for irrigation and suction using the Patient Side Manipulator (PSM) from the dVRK and the EndoWrist Suction/Irrigator. Screen-space fluid rendering is implemented and adapted to simulate realistic visual effects, including fluid mixing.
  • With hand-crafted reward functions, DR, curriculum learning (CL), and IL, two vision-based agents for irrigation and suction are trained in the simulator for controlling the robot in joint space to complete the tasks.
@@ II-A Surgical Subtask Automation and Augmented Dexterity
A large number of recent studies aim to achieve subtask automation in surgeries using surgical robotic systems, with a long-term goal of increasing the level of autonomy in surgeries, similar to autonomous vehicles. The concept where surgical subtasks are controlled by a robot under surgeons’ supervision is recently referred to as augmented dexterity [11]. Examples of these studies include tissue cutting [1, 2, 3], tissue manipulation [5, 6, 7], suturing [8, 9, 10], blood suction [16, 17, 18, 19, 20], and vessel manipulation [23]. Numerous studies have also focused on the navigation and control of endoscopes [24, 25, 26]. Many of these studies leverage RL approaches [1, 2, 3, 5, 6, 7, 9, 10, 8, 19]. However, applying RL to surgical tasks involving fluid manipulation, such as irrigation and suction, remains relatively underexplored due to the challenges and limited software tools for simulating fluid dynamics.
@@ II-B Surgical Simulation and Sim-to-Real Transfer
Surgical task simulation is challenging compared with daily tasks, due to the multiple types of objects and manipulations involved, such as soft bodies (e.g., soft tissue), fluids (e.g., blood and other body fluids), burning, and cutting. Nevertheless, its applications have been important in two major domains. The traditional usage is for training surgeons, where trainees operate in a simulated surgical environment to improve their skills. Examples of such systems include da Vinci SimNow111www.intuitive.com/en-us/products-and-services/da-vinci/learning/simnow, VirtaMed LaproS222www.virtamed.com/en/products-and-solutions/simulators/laparos, and iMSTK333www.imstk.org.
A more emerging application of surgical simulation is for training autonomous agents using robot learning methods, where agents are trained in the simulator and transferred to the real world. Studies of this type include [4, 7, 9, 10, 13, 19]. While it is ideal to use commercial simulators such as the da Vinci SimNow for building training environments for robot learning, it is challenging in practice due to the proprietary nature of these software systems. Many recent studies have proposed open-source simulation environments for surgical robot learning, including dVRL [27], AMBF-RL [28], UnityFlexML [4], SurRoL [29, 30], LapGym [31], Surgical Gym [32], and ORBIT-Surgical [33]. Among them, a number of recent studies consider GPU-accelerated simulation, which allows large-scale parallel training and significantly boosts training efficiency and performance. However, they are generally inferior to the commercial ones in terms of the types of objects that they can simulate. For instance, AMBF-RL is mainly used for simulating rigid bodies and some soft bodies, without the capability for simulating fluids. Nevertheless, achieving sim-to-real transfer for irrigation-suction requires the simulation of visually plausible fluids. There are also studies that build task-specific simulation environments, such as [12, 13].
@@ II-C Fluid Simulation and Manipulation
This work is also related to recent advancements in robotic manipulation involving fluids, as well as the development of fluid simulation environments for robot learning [34, 35, 36]. In [34], the Navier–Stokes equation is used for simulating 2D fluid flow, and an RL model is trained to learn to manipulate a rigid object using a water jet. Babaians et al. used PBF for simulating fluids [35], which is used to train a robot to pour a glass of liquid. Authors of [36] considered a number of fluid manipulation tasks, such as fluid mixing and drawing latte art. Most of these studies focus on daily tasks such as pouring and beverage mixing, which is different from surgical irrigation or suction in terms of manipulation contact. For instance, daily tasks rarely involve fluid suction that requires the simulation of a suction force.
There are a number of real-time fluid simulation and animation techniques, such as smoothed-particle hydrodynamics (SPH), material point method (MPM), fluid implicit particle (FLIP), and PBF [37]. Among these methods, PBF is one of the most computationally efficient and numerically stable ones due to its consideration of positional constraints instead of relying on force integration. This makes it particularly well-suited for large-scale trial-and-error learning, enabling massive parallel training, such as in [35], while avoiding pauses or time rewinds caused by numerical instability. For fluid rendering, there are two main technical approaches: (a) mesh-based surface reconstruction, and (b) screen-space rendering. The first approach reconstructs a surface mesh from fluid particles or level sets using techniques such as marching cubes [38]. However, this approach is generally more computationally expensive than the screen-space method, which approximates the fluid surface directly in image space using depth maps or normal reconstruction techniques [39, 40, 41]. This is widely used in real-time applications, often in combination with PBF to achieve visually plausible results with less computation.
@@ III-A1 Fluid Simulation With Color Diffusion
To simulate irrigation and suction, fluid behavior must be simulated and rendered with visual plausibility. PhysX 5 SDK provides the feature of simulating fluid particles on GPU using position-based dynamics (PBD), and more specifically, position-based fluids (PBF) [42]. Although PhysX 5 includes an existing PBF implementation, it focuses solely on numerical computation without rendering considerations. In this work, we further introduce a straightforward approach to fluid color mixing when using PhysX’s PBF.
@@
As we would like to mix two types of fluids with different colors when simulating irrigation, the simulation should include color mixing and changes. It is therefore straightforward to label each particle with a color and diffuse the colors between particles based on their spatial proximity. To achieve this, each particle is assigned to a cell in a 3D grid with uniform discrete cells based on its position, also known as spatial hashing. For each particle, all particles in a neighboring region are evaluated, and the weighted average of the particle colors is computed to update the particle’s color. Therefore, for each particle $i$ with color $\mathbf{c}_{i}$,
@@ III-A2 Fluid Rendering
A sequence of screenshots from a demonstration scene illustrating the simulation’s capability in fluid physics and rendering. The images showcase the interaction between two flowing fluids, their mixing behavior, and visual effects.
@@
However, rendering should account for the different colors of each particle in our case. To achieve this, an additional color extraction pass is added to the regular screen-space rendering scheme after the surface depth texture is extracted, which simply extracts the colors of particles that are closest to the fluid surface. This process is similar to [43], where the authors propose using a separate pass to extract the colors based on the thickness of the fluid. The difference is that we do not consider any particle colors apart from the surface ones. This eliminates the need for a thickness pass to extract the fluid thickness and is a valid simplification without losing much visual plausibility, thanks to the previously discussed color diffusion scheme.
We use anisotropy ellipsoids discussed in [41] when reconstructing the fluid surface and a narrow range filter [44] to smooth it. An example fluid simulation and rendering scene is shown in Fig. 3d. The overall fluid rendering steps are shown in Fig. 3. In a stress test scene with two large mixing fluid blocks of different colors, each containing 125 thousand particles, the average render time is approximately 7.5 ms per frame (measured over multiple frames). The physics computation is the primary bottleneck, requiring an average of 24.1 ms per physics step. Considering task parallelism and all other operations, the frame rate can generally be achieved around 30 Hz. This test was conducted with an Nvidia RTX 4070, a mid-tier consumer GPU. As a comparison, a scene with 20 robots takes less than 3 ms for physics computation per step.
@@ III-B A Surgical Robot Learning Framework for the dVRK
In our previous work [22], a general surgical simulation platform (CRESSim) has been built by leveraging PhysX 5 and Unity, where the dVRK PSM robot is simulated and the real-world dVRK MTM robot is used for teleoperating the simulated PSM. The PSM is a robot with a mechanical remote center of motion (RCM). The robot, together with the EndoWrist One Suction/Irrigator, is simulated as a kinematic tree with multiple articulation joints, as shown in Fig. 4. It is worth noting that the Suction/Irrigator has only two degrees of freedom (DoF), making the entire PSM a 5-DoF robot.
@@
Based on this, we are able to further implement a robot learning framework by incorporating the Unity Machine Learning Agents Toolkit (ML-Agents) [45] into the existing software. To allow parallel training environments, additional modifications were made for long-term episodic resetting of the environments, such as fluid particle and material property resetting. Additional features are introduced, such as shared object meshes and shapes between duplicate training areas, which can potentially save memory usage for large-scale training. Human demonstrations for completing simulated tasks can then be collected while using the real-world MTM to teleoperate the PSM in the simulated learning scene, as shown in Fig 5. The procedures for collecting demonstrations will be further discussed in Section IV-B.
Figure 5: A surgical robot learning platform for the dVRK with Unity, ML-Agents, and PhysX 5. The human operator sees the simulated camera in the console and teleoperates the simulated robot for collection demonstrations.
Based on this platform, two simulated learning scenes for irrigation and suction are developed, which will be discussed in the following section. While we only consider irrigation and suction in this work, the platform can be generally used for building various surgical robot learning tasks for the dVRK, including those that use different instruments such as the Large Needle Driver. Previous work has shown its capability to simulate a number of surgical tasks, such as tissue cutting and manipulation [22]. As the proposed platform combines CRESSim with ML-Agents, we refer to it as CRESSim-ML.